# Cosmic Collider Gravitational Waves sourced by Right-handed Neutrino production from Bubbles: Testing Seesaw, Leptogenesis and Dark Matter

### v1

@@ Abstract
We study a minimal type-I seesaw framework in which a first-order phase transition (FOPT), driven by a singlet scalar $\phi$, produces right-handed neutrinos (RHNs) through bubble collisions, realizing a cosmic-scale collider that probes ultra-high energy scales. The resulting inhomogeneous RHN distribution sources a novel low-frequency gravitational-wave (GW) signal in addition to the standard bubble-collision contribution. A stable lightest RHN can account for the observed dark matter (DM) relic abundance for masses as low as $M_{1}\equiv m_{\rm DM}\gtrsim 10^{6},\mathrm{GeV}$, with the associated novel GW signal accessible in LISA, ET and upcoming LVK detectors. If the RHNs are unstable, their CP-violating decays generate the baryon asymmetry via leptogenesis for $M_{1}\gtrsim 10^{11},\mathrm{GeV}$ and phase transition temperatures $T_{*}\gtrsim 10^{6}\,\mathrm{GeV}$, yielding the novel GW signatures within the reach of ET, BBO and upcoming LVK detectors. Part of the parameter space is already constrained by the LVK $\mathcal{O}(3)$ data. If RHN decays populate a dark sector fermion $\chi$ with mass $m_{\chi}\in[10^{-4},10^{4}],\mathrm{GeV}$, successful co-genesis of baryons and asymmetric dark matter is achieved for $T_{*}\gtrsim 10^{7}\,\mathrm{GeV}$ and $M_{1}\gtrsim 10^{9}\,\mathrm{GeV}$, naturally explaining $\Omega_{\rm DM}\simeq 5\Omega_{\rm B}$. The corresponding GW signals are testable with LISA, ET, and BBO. Finally, we analyze a UV-complete multi-Majoron model based on a global $U(1)_{N}\times U(1)_{\rm B-L}$ extension, in which a distinctive GW signature associated with a cosmic Majoron collider arises from $\varphi_{1}$ production during $U(1)_{N}$ symmetry breaking, detectable by BBO, ET and upcoming LVK. Successful leptogenesis is realized for $M_{3}\sim 10^{10}\,\mathrm{GeV}$ and a $U(1)_{N}$ breaking vev $v_{2}\sim\mathcal{O}(\mathrm{TeV})$ that sets the seesaw scale.
@@ I Introduction
Cosmological first-order phase transitions (FOPTs) involve the nucleation of true-vacuum bubbles in a metastable false-vacuum background Hogan:1983ixn; Witten:1984rs; Hogan:1986qda; Kosowsky:1991ua; Kamionkowski:1993fg; Kosowsky:1992vn. When these bubbles expand and collide, they generate a large stochastic gravitational-wave (GW) background that can be tested in upcoming GW missions Grojean:2006bp; Caprini:2015zlo; Caprini:2018mtu; Caprini:2019egz; Athron:2023xlk.
@@
In certain classes of FOPTs, the bubble walls may accelerate to ultra-relativistic speeds, the so-called runaway scenario. This typically occurs during supercooled transitions, transitions into cold sectors, or sectors lacking gauge bosons. In such cases, plasma friction becomes negligible due to plasma dilution, and a large fraction of the released vacuum energy accumulates in the bubble walls. Consequently, the GW signal is dominated by the scalar field energy in the walls after collision Kosowsky:1991ua; Kosowsky:1992rz; Kosowsky:1992vn; Kamionkowski:1993fg; Caprini:2007xq; Huber:2008hg; Bodeker:2009qy; Jinno:2016vai; Jinno:2017fby; Konstandin:2017sat; Cutting:2018tjt; Cutting:2020nla; Lewicki:2020azd.
@@
Beyond the classical picture of bubble dynamics and energy dissipation Hawking:1982ga; Kleban:2011pg; Chang:2008gj, ultra-relativistic walls can also produce particles with energies far above the ambient plasma temperature, an inherently quantum effect analogous to particle production from vacuum. Collisions of such vacuum bubbles in the early Universe thus serve as as cosmic-scale high-energy colliders with energy even upto the Planck scale. These “cosmic colliders” [^1] Not to be confused with ”Cosmological Collider” which involves particle production during inflation from inflationary fluctuations Arkani-Hamed:2015bza. are then possibly the most energetic phenomena in our reached ever Shakya:2025qpi, or larger than any energy scale reachable in traditional laboratory colliders. On microscopic scales (set by the Lorentz-contracted bubble-wall thickness), bubble collisions can efficiently convert vacuum energy into energetic particles Jinno:2017fby; Konstandin:2011ds; Falkowski:2012fb; Watkins:1991zt. Recent work Shakya:2023kjf; Mansour:2023fwj shows that this production follows a universal power law that is largely insensitive to the details of the collision, provided the walls are ultra-relativistic. If the produced particles survive long enough to generate a quadrupole anisotropy, they can source an additional GW component Inomata:2024rkt. Such GWs can probe fundamental particle-physics properties and offer new windows into early-Universe dynamics.
@@
Observations of neutrino oscillations Fukuda_2001; Fukuda_2002; Ahmad_2002; Ashie_2005; Abe_2016; Smirnov_2016; Aartsen_2018; Abe_2008; Abe_2011 imply that neutrinos possess non-zero masses, requiring physics beyond the Standard Model (SM). Introducing two or more right-handed Majorana neutrinos (RHNs) leads to the Type-I seesaw mechanism Minkowski:1977sc; Yanagida:1979as; Glashow:1979nm; Mohapatra:1979ia, which naturally explains small neutrino masses. The same RHNs can also account for the observed baryon asymmetry of the Universe (BAU) Planck2018; ParticleDataGroup:2020ssz through the leptogenesis mechanism Fukugita:1986hr. In conventional thermal leptogenesis, successful asymmetry generation requires the lightest RHN to satisfy the Davidson–Ibarra bound, $M_{1}\gtrsim 10^{9}$ GeV, assuming the RHN mass spectrum is hierarchical, along with no lepton flavor effects Nardi:2006fx; Abada:2006fw; Abada:2006ea; Giudice:2003jh. Such high scales challenge direct tests in laboratory experiments, however, certain indirect signatures of new physics like lepton number violation processes through neutrinoless double beta decay Cirigliano:2022oqy or CP violation in neutrino oscillation Endoh:2002wm can be looked for. Moreover, theoretical constraints on low-energy couplings can arise from demanding consistency with UV-complete frameworks, such as $SO(10)$ Grand Unified Theories (GUTs) DiBari:2008mp; Bertuzzo:2010et; Buccella:2012kc; Altarelli:2013aqa; Fong:2014gea; Mummidi:2021anm; Patel:2022xxu. There may be additional consistencies related to the demand of electroweak (EW) SM Higgs vacuum (meta)stability in the early Universe Ipek:2018sai; Croon:2019dfw. All of these provide some useful bounds on the large parameter space involved in Seesaw and Leptogenesis scenarios. To further make these bounds sharper and probe the relevant high scales of leptogenesis, we allude to non traditional or non laboratory based searches searches, particularly cosmological observables as tools to see indirect signatures or tests for such scales of new physics involving seesaw, leptogenesis, and heavy dark matter etc.
@@
The other major puzzle is the nature and origin of dark matter (DM), which constitutes about $\sim 27\%$ of the Universe’s energy density Zwicky:1933gu; Zwicky:1937zza; Rubin:1970zza; Clowe:2006eq. In terms of density parameter $\Omega_{\rm DM}$ and reduced Hubble constant $h=\text{Hubble Parameter}/(100\;\text{km}~\text{s}^{-1}\text{Mpc}^{-1})$, the current DM abundance is usually depicted to be Aghanim:2018eyx
@@
Figure 1: Schematic diagram for Cosmic Collider Shakya:2025qpi during runaway FOPTs showing the correlation between RHN production, Dark matter and Leptogenesis along with the resulting cosmological signals - the well established bubble collision GWs as well as the novel RHN production GWs as discussed in this analysis.
The observed BAU is quantitatively depicted as the ratio of excess of baryons over anti-baryons to photon Aghanim:2018eyx
@@
as observed from the cosmic microwave background (CMB) measurements, and from the big bang nucleosynthesis (BBN) estimates  Zyla:2020zbs. The numerical proximity of DM and baryon abundances $\Omega_{\rm DM}\approx 5\,\Omega_{\rm Baryon}$, motivates scenarios where both originate from a common mechanism (see Ref. Boucenna:2013wba for a review on this topic). This has led to many proposals where the usual mechanism for baryogenesis directly or via leptogenesis is extended to the dark sector assuming the dark sector to be asymmetric Nussinov:1985xr; Davoudiasl:2012uw; Petraki:2013wwa; Zurek:2013wia. In typical asymmetric dark matter (AsDM) scenario, the same out-of-equilibrium decay of a heavy particle, for instance, the decay of heavy right-handed neutrinos (RHN) into SM leptons and dark sector leads to the creation of asymmetries in the two sectors of similar order of magnitudes $n_{B}-n_{\overline{B}}\sim\lvert n_{\rm DM}-n_{\overline{\rm DM}}\rvert$ Falkowski_2011; Barman_2022.
In order to understand the impact of the presence of such heavy neutrinos in the theory, we investigate FOPT in a minimal scalar singlet ($\phi$) extended SM. During the phase transition, the bubble walls of $\phi$ decays to RHNs. The subsequent decays of these right-handed neutrinos are responsible for baryogenesis via leptogenesis Fukugita:1986hr. We also study a simplistic possibility where one of the right-handed neutrino is stable and be the DM candidate. Finally we also investigate the scenario where RHN decays can also furnish a dark matter production mechanism Falkowski_2011; Barman_2022, which can result in asymmetric dark matter.
@@ II Particle Production from Bubbles during first-order Phase Transition
During a FOPT bubble wall collisions can lead to production of various particles coupled to the scalar field . Depending on the details of the scalar potential, the collisions can be elastic or inelastic. In both cases, the scalar field at the point of collision gets excited to a field value away from the minima. This results in oscillations around true (false) minimum for inelastic (elastic) collisions. Such dynamics of the background field then can give rise to significant particle production to any field that couples to it leading to a significant fraction of the vacuum energy getting converted into particle population cataldi2024leptogenesisbubblecollisions; shakya2025aspectsparticleproductionbubble; Watkins:1991zt.
@@ II.1 Formalism
The formalism consists of treating the moving and colliding bubble walls and subsequent oscillations as classical external field configurations of the scalar $\phi(x,t)$. The probability for this configuration to decay into particles is extracted from the imaginary part of its effective action Watkins:1991zt,
@@ II.2 The Efficiency Factor
The efficiency factor $f(p^{2})$ depends on the nature of the collision and the subsequent dynamics of the background field and has been calculated in the literature. We quote the results here for a runaway phase transition with negligible plasma friction, which will be used throughout this work Mansour:2023fwj:
@@
We see this has an approximately power law component $f_{PE}\sim p^{-4}$, which originates from the nontrivial dynamics of the background field when the bubbles collide. The remaining part in $f(p^{2})$ behaves approximately as a Gaussian peak centered around the mass of the scalar in the relevant vacuum. This comes from the oscillation of the scalar field around its relevant minimum after the collision. We note shakya2025aspectsparticleproductionbubble from numerical estimates, the log factor in Eq. (15) ranges between 6 $\sim$ 60 for a wide range of phase transition parameter values and hence can be treated as a constant.
@@ II.3 Effective Action Calculation
    Fermions: Fermions can be produced via some Yukawa coupling to the scalar $\phi$ of the form $\mathcal{L}_{I}\supset-y_{f}\phi\bar{\psi}\psi$. This gives rise to $\phi^{*}\rightarrow\bar{\psi}\psi$ decay process Mansour:2023fwj:
@@ III Gravitational Waves from FOPT during Particle Production
Gravitational waves (GWs) from first-order phase transitions (FOPTs) in the early Universe Witten:1984rs; Hogan:1986qda; PhysRevLett.65.3080; Kamionkowski_1994 have been widely studied as well-motivated and promising targets for current and upcoming GW detectors amaroseoane2017laserinterferometerspaceantenna; Kawamura:2006up; Punturo:2010zz. Although the electroweak phase transition and the QCD phase transition within the Standard Model (SM) are known to not be first-order [^2] Though some new possibilities has been recently proposed in literature, see Ref. Shakya:2025mdh, the detection of a cosmological FOPT would be a smoking gun for physics beyond the Standard Model (BSM). Indeed, there are numerous BSM frameworks that predict FOPTs in dark sectors Schwaller_2015; Caprini_2018; Baldes_2017; Hall_2023; Croon_2018; Prokopec_2019 which can produce observable GW signals.
@@ III.1 Theoretical Framework for Gravitational Waves
then we find the Gravitational wave energy densityCaprini_2018

### v2

# Cosmic Collider Gravitational Waves sourced by Right-handed Neutrino production from Bubbles: Testing Scales of Seesaw, Leptogenesis and Dark Matter

@@ Abstract
We study a minimal type-I seesaw framework in which a first-order phase transition (FOPT), driven by a singlet scalar $\phi$, produces right-handed neutrinos (RHNs) through bubble collisions, realizing a cosmic-scale collider that probes ultra-high energy scales. The resulting inhomogeneous RHN distribution sources a novel low-frequency gravitational-wave (GW) signal in addition to the standard bubble-collision contribution. A stable lightest RHN can account for the observed dark matter (DM) relic abundance for masses as low as $M_{1}\equiv m_{\rm DM}\gtrsim 10^{6}\,\mathrm{GeV}$, with the associated novel GW signal accessible in LISA, ET and upcoming LVK detectors. If the RHNs are unstable, their CP-violating decays generate the baryon asymmetry via leptogenesis for $M_{1}\gtrsim 10^{11}\,\mathrm{GeV}$ and phase transition temperatures $T_{*}\gtrsim 10^{6}\,\mathrm{GeV}$, yielding the novel GW signatures within the reach of ET, BBO and upcoming LVK detectors. Part of the parameter space is already constrained by the LVK $\mathcal{O}(3)$ data. If RHN decays populate a dark sector fermion $\chi$ with mass $m_{\chi}\in[10^{-4},10^{4}],\mathrm{GeV}$, successful co-genesis of baryons and asymmetric dark matter is achieved for $T_{*}\gtrsim 10^{7}\,\mathrm{GeV}$ and $M_{1}\gtrsim 10^{9}\,\mathrm{GeV}$, naturally explaining $\Omega_{\rm DM}\simeq 5\Omega_{\rm B}$. The corresponding GW signals are testable with LISA, ET, and BBO. Finally, we analyze a UV-complete multi-Majoron model based on a global $U(1)_{N}\times U(1)_{\rm B-L}$ extension of Standard Model (SM), motivated from the hierarchy of lepton masses, in which a distinctive GW signature associated with a cosmic Majoron collider arises from scalar production during $U(1)_{N}$ symmetry breaking, detectable by BBO, ET and upcoming LVK. Successful leptogenesis is realized for the heaviest RHN mass $M_{3}\sim 10^{10}\,\mathrm{GeV}$ and a $U(1)_{N}$ breaking vev $v_{2}\sim\mathcal{O}(\mathrm{TeV})$ that sets the seesaw scale.
@@ I Introduction
Cosmological first-order phase transitions (FOPTs) involve the nucleation of true-vacuum bubbles in a metastable false-vacuum background Hogan (1983); Witten (1984); Hogan (1986); Kosowsky et al. (1992a); Kamionkowski et al. (1994a); Kosowsky and Turner (1993). When these bubbles expand and collide, they generate a large stochastic gravitational-wave (GW) background that can be tested in upcoming GW missions Grojean and Servant (2007); Caprini and others (2016); Caprini and Figueroa (2018a); Caprini and others (2020); Athron et al. (2024).
@@
In certain classes of FOPTs, the bubble walls may accelerate to ultra-relativistic speeds, the so-called runaway scenario. This typically occurs during supercooled transitions, transitions into cold sectors, or sectors lacking gauge bosons. In such cases, plasma friction becomes negligible due to plasma dilution, and a large fraction of the released vacuum energy accumulates in the bubble walls. Consequently, the GW signal is dominated by the scalar field energy in the walls after collision Kosowsky et al. (1992a, b); Kosowsky and Turner (1993); Kamionkowski et al. (1994a); Caprini et al. (2008); Huber and Konstandin (2008); Bodeker and Moore (2009); Jinno and Takimoto (2017b, 2019b); Konstandin (2018); Cutting et al. (2018b, 2021); Lewicki and Vaskonen (2021).
@@
Beyond the classical picture of bubble dynamics and energy dissipation Hawking et al. (1982); Kleban (2011); Chang et al. (2009), ultra-relativistic walls can also produce particles with energies far above the ambient plasma temperature, an inherently quantum effect analogous to particle production from vacuum. Collisions of such vacuum bubbles in the early Universe thus serve as as cosmic-scale high-energy colliders with energy even upto the Planck scale. These “cosmic colliders” [^1] Not to be confused with “Cosmological Collider” which involves particle production during inflation from inflationary fluctuations Arkani-Hamed and Maldacena (2015). are then possibly the most energetic phenomena in our universe reached ever Shakya (2025a), and have energies larger than any energy scale reachable in traditional laboratory colliders. On microscopic scales (set by the Lorentz-contracted bubble-wall thickness), bubble collisions can efficiently convert vacuum energy into energetic particles Jinno and Takimoto (2019b); Konstandin and Servant (2011); Falkowski and No (2013); Watkins and Widrow (1992). Recent work Shakya (2025b); Mansour and Shakya (2025) shows that this production follows a universal power law that is largely insensitive to the details of the collision, provided the walls are ultra-relativistic. If the produced particles survive long enough to generate a quadrupole anisotropy, they can source an additional GW component Inomata et al. (2025). Such GWs can probe fundamental particle-physics properties and offer new windows into early-Universe dynamics.
@@
Observations of neutrino oscillations Fukuda and others (2001, 2002); Ahmad and others (2002); Ashie and others (2005); Abe and others (2016); Smirnov and others (2016); Aartsen and others (2018); Abe and others (2008, 2011) imply that neutrinos possess non-zero masses, requiring physics beyond the Standard Model (SM). Introducing two or more right-handed Majorana neutrinos (RHNs) leads to the Type-I seesaw mechanism Minkowski (1977); Yanagida (1979); Glashow (1980); Mohapatra and Senjanovic (1980), which naturally explains small neutrino masses. The same RHNs can also account for the observed baryon asymmetry of the Universe (BAU) Collaboration (2020); Zyla and others (2020a) through the leptogenesis mechanism Fukugita and Yanagida (1986). In conventional thermal leptogenesis, successful asymmetry generation requires the lightest RHN to satisfy the Davidson–Ibarra bound, $M_{1}\gtrsim 10^{9}$ GeV, assuming the RHN mass spectrum is hierarchical, along with no lepton flavor effects Nardi et al. (2006); Abada et al. (2006b, a); Giudice et al. (2004). Such high scales challenge direct tests in laboratory experiments, however, certain indirect signatures of new physics like lepton number violation processes through neutrinoless double beta decay Cirigliano and others (2022) or CP violation in neutrino oscillation Endoh et al. (2002) can be looked for. Moreover, theoretical constraints on low-energy couplings can arise from demanding consistency with UV-complete frameworks, such as $SO(10)$ Grand Unified Theories (GUTs) Di Bari and Riotto (2009); Bertuzzo et al. (2011); Buccella et al. (2012); Altarelli and Meloni (2013); Fong et al. (2015); Mummidi and Patel (2021); Patel (2023). There may be additional consistencies related to the demand of electroweak (EW) SM Higgs vacuum (meta)stability in the early Universe Ipek et al. (2018); Croon et al. (2019). All of these provide some useful bounds on the large parameter space involved in Seesaw and Leptogenesis scenarios. To further make these bounds sharper and probe the relevant high scales of leptogenesis, we allude to non traditional or non laboratory based searches, particularly cosmological observables as tools to see indirect signatures or tests for such scales of new physics involving seesaw, leptogenesis, and heavy dark matter etc.
@@
The other major puzzle is the nature and origin of dark matter (DM), which constitutes about $\sim 27\%$ of the Universe’s energy density Zwicky (1933, 1937); Rubin and Ford (1970); Clowe et al. (2006). In terms of density parameter $\Omega_{\rm DM}$ and reduced Hubble constant $h=\text{Hubble Parameter}/(100\;\text{km}~\text{s}^{-1}\text{Mpc}^{-1})$, the current DM abundance is usually depicted to be Aghanim and others (2020a)
@@
Figure 1: Schematic diagram for Cosmic Collider Shakya (2025a) during runaway FOPTs showing the correlation between RHN production, Dark matter and Leptogenesis along with the resulting cosmological signals - the well established bubble collision GWs as well as the novel RHN production GWs as discussed in this analysis.
The observed BAU is quantitatively depicted as the ratio of excess of baryons over anti-baryons to photon Aghanim and others (2020a)
@@
as observed from the cosmic microwave background (CMB) measurements, and from the big bang nucleosynthesis (BBN) estimates  Zyla and others (2020b). The numerical proximity of DM and baryon abundances $\Omega_{\rm DM}\approx 5\,\Omega_{\rm Baryon}$, motivates scenarios where both originate from a common mechanism (see Ref. Boucenna and Morisi (2014) for a review on this topic). This has led to many proposals where the usual mechanism for baryogenesis directly or via leptogenesis is extended to the dark sector assuming the dark sector to be asymmetric Nussinov (1985); Davoudiasl and Mohapatra (2012); Petraki and Volkas (2013); Zurek (2014). In typical asymmetric dark matter (AsDM) scenario, the same out-of-equilibrium decay of a heavy particle, for instance, the decay of heavy right-handed neutrinos (RHN) into SM leptons and dark sector leads to the creation of asymmetries in the two sectors of similar order of magnitudes $n_{B}-n_{\overline{B}}\sim\lvert n_{\rm DM}-n_{\overline{\rm DM}}\rvert$ Falkowski et al. (2011a); Barman et al. (2022).
In order to understand the impact of the presence of such heavy neutrinos in the theory, we investigate FOPT in a minimal scalar singlet ($\phi$) extended SM. During the phase transition, the bubble walls of $\phi$ decays to RHNs. The subsequent decays of these right-handed neutrinos are responsible for baryogenesis via leptogenesis Fukugita and Yanagida (1986). We also study a simplistic possibility where one of the right-handed neutrino is stable and be the DM candidate. Finally we also investigate the scenario where RHN decays can also furnish a dark matter production mechanism Falkowski et al. (2011a); Barman et al. (2022), which can result in asymmetric dark matter.
@@ II Particle Production from Bubbles during first-order Phase Transition
During a FOPT bubble wall collisions can lead to production of various particles coupled to the scalar field . Depending on the details of the scalar potential, the collisions can be elastic or inelastic. In both cases, the scalar field at the point of collision gets excited to a field value away from the minima. This results in oscillations around true (false) minimum for inelastic (elastic) collisions. Such dynamics of the background field then can give rise to significant particle production to any field that couples to it leading to a significant fraction of the vacuum energy getting converted into particle population Cataldi and Shakya (2024); Shakya (2025c); Watkins and Widrow (1992).
@@ II.1 Formalism
The formalism consists of treating the moving and colliding bubble walls and subsequent oscillations as classical external field configurations of the scalar $\phi(x,t)$. The probability for this configuration to decay into particles is extracted from the imaginary part of its effective action Watkins and Widrow (1992),
@@ II.2 The Efficiency Factor
The efficiency factor $f(p^{2})$ depends on the nature of the collision and the subsequent dynamics of the background field and has been calculated in the literature. We quote the results here for a runaway phase transition with negligible plasma friction, which will be used throughout this work Mansour and Shakya (2025):
@@
We see this has an approximately power law component $f_{PE}\sim p^{-4}$, which originates from the nontrivial dynamics of the background field when the bubbles collide. The remaining part in $f(p^{2})$ behaves approximately as a Gaussian peak centered around the mass of the scalar in the relevant vacuum. This comes from the oscillation of the scalar field around its relevant minimum after the collision. We note Shakya (2025c) from numerical estimates, the log factor in Eq. (15) ranges between 6 $\sim$ 60 for a wide range of phase transition parameter values and hence can be treated as a constant.
@@ II.3 Effective Action Calculation
    Fermions: Fermions can be produced via some Yukawa coupling to the scalar $\phi$ of the form $\mathcal{L}_{I}\supset-y_{f}\phi\bar{\psi}\psi$. This gives rise to $\phi^{*}\rightarrow\bar{\psi}\psi$ decay process Mansour and Shakya (2025):
@@ III Gravitational Waves from FOPT during Particle Production
Gravitational waves (GWs) from first-order phase transitions (FOPTs) in the early Universe Witten (1984); Hogan (1986); Turner and Wilczek (1990); Kamionkowski et al. (1994b) have been widely studied as well-motivated and promising targets for current and upcoming GW detectors Amaro-Seoane et al. (2017); Kawamura and others (2006); Punturo and others (2010). Although the electroweak phase transition and the QCD phase transition within the Standard Model (SM) are known to not be first-order [^2] Though some new possibilities has been recently proposed in literature, see Ref. Shakya (2025d); Yamada (2025, 2026)., the detection of a cosmological FOPT would be a smoking gun for physics beyond the Standard Model (BSM). Indeed, there are numerous BSM frameworks that predict FOPTs in dark sectors Schwaller (2015); Caprini and Figueroa (2018b); Baldes (2017); Hall et al. (2023); Croon et al. (2018); Prokopec et al. (2019) which can produce observable GW signals.
@@ III.1 Theoretical Framework for Gravitational Waves
then we find the Gravitational wave energy densityCaprini and Figueroa (2018b)